Write the expression as a sum and or a difference of logarithms with all variables to the first degree.
ln sqrt(fz)divided by z^10
Answer: 1/2 lnf - 19/2 lnz

Answer by stanbon(57285)

Write the expression as a sum and or a difference of logarithms with all variables to the first degree.
ln sqrt(fz)divided by z^10
----
ln[sqrt[fz/z^10]
---
= (1/2)ln[fz/z^10]
= (1/2[ln[f/z^9]
= (1/2)[ln(f)-9ln(z)]
= (1/2)ln(f) - (9/2)ln(z)
================================
Cheers,
Stan H.